Electric Potential

A uniform electric field of magnitude 5.1×105 points in the positive direction.

a.)Find the change in electric potential between the origin and the point (0, 6.0 ).
b.)Find the change in electric potential between the origin and the point (6.0 , 0).
c.)Find the change in electric potential between the origin and the point (6.0 , 6.0 ).
Answer #1
Electric field, E = 5.1 x 10^5 V/m(a)Distance, d = 6 mChange in electric potential = E d= 5.1 x 10^5 * 6= 30.6 x 10^5 V(b)Distance, d = 6 mChange in electric potential = E d= 5.1 x 10^5 * 6= 30.6 x 10^5 V(c)(a)Distance, d = √(6^2 + 6^2) m = 8.49 mChange in electric potential = E d= 5.1 x 10^5 * 8.49=4.33 x 10^6 V
answered by: Moncha